# When Respondents Don’t Care Anymore: Identifying the Onset of Careless Responding

### v3

@@ Results
Third, there may be participants who alternate between periods of attentive and careless responding (Berry et al.,, 1992; Clark et al.,, 2003; Meade & Craig,, 2012). Our simulations indicate that CODERS remains conservative in this case and rarely flags one of the changepoints (see Appendix C.5). CODERS should therefore be viewed as a promising first step towards detecting periods of careless responding, with an extension for the detection of multiple changepoints being an area of further research.
Finally, the changepoint detection procedure of Shao & Zhang, (2010) employed by CODERS is based on asymptotic arguments when the number of items is large. Hence, the number of items in the questionnaire is the main determinant for the statistical power of CODERS. For the detection of inconsistent careless responding, the number of participants plays a secondary role regarding power: the more attentive responses the autoencoder can learn from, the better the reconstruction errors can distinguish between attentive and careless responding (cf. Appendix C.3). The LSP sequence of a given participant, on the other hand, depends only on their own responses, meaning that the number of observations should not play a role concerning power in detecting invariable careless responding. CODERS is thus primarily designed for relatively lengthy questionnaires, and our simulations suggest that there is a gradual drop in statistical power for detecting inconsistency as survey length decreases (see Appendix C.2). However, partial carelessness due to survey fatigue may be less prevalent in short questionnaires (cf. Bowling et al.,, 2021). A detailed study of partially careless responding in shorter surveys is therefore beyond the scope of this paper.
@@
In addition to inconsistency and invariability, fast responding is considered the third major manifestation of carelessness (e.g., Ward & Meade,, 2023). In (online) surveys, response time is typically measured by the total time a participant spent on the questionnaire, time spent on each questionnaire page, or (less common) time spent on each questionnaire item.
If per-item response times are available, they can easily be incorporated into CODERS as an additional series to be used in the changepoint detection. If only per-page response times are available and pages are relatively short, an additional series can be obtained by assigning to each response the average page time, i.e., the time the participant spent on the page on which the item is located divided by the number of items on that page (which is inspired by the page time index of Bowling et al.,, 2023). If a participant at some point in the survey, for instance, starts responding without reading the items, we expect a changepoint in response time towards considerably faster responses (cf. Bowling et al.,, 2023, 2016; Meade & Craig,, 2012; Huang et al.,, 2012).
Nevertheless, it is unclear if a changepoint in response times is indeed indicative of careless responding in an empirical setting. This may depend on the survey design: if items in one part of the survey are sufficiently longer than in other parts, a changepoint in response times might occur naturally for attentive respondents. On the other hand, if all items are of similar length, a changepoint to (much) shorter response times may indicate carelessness. For instance, there exist instruments whose individual items comprise of single words, such as the Big Five unipolar markers (Goldberg,, 1992) with items like “timid” or “envious”. In such measurements, one would not expect large differences in response times across items. For such survey designs with similar item length, incorporating response times into CODERS may prove fruitful, but a thorough validation is beyond the scope of this paper and therefore left for future research.
@@ Discussion
We introduced a novel method, CODERS, designed to identify the onset of careless responding (or absence thereof) for each participant in a survey. It is motivated by literature stressing that survey length contributes to careless responding, in the sense that lengthy surveys—which are common in the behavioral and organizational sciences—may result in a large proportion of survey participants being partially careless, for instance due to fatigue or boredom. CODERS is highly flexible and can combine multiple indicators for the potential onset of carelessness, such as autoencoder reconstruction errors (RE) as an indicator of inconsistency, and the proposed LongStringPattern (LSP) sequence as an indicator of invariability. Simulation experiments indicate excellent performance of CODERS in identifying the onset of careless responding, and highlight the importance of incorporating multiple indicators that capture different manifestations of carelessness. Re-analyzing data from a seminal study on careless responding, we demonstrate how CODERS reveals novel insights on partial carelessness. In particular when combined with a post-hoc analysis of the identified response segments, CODERS is able to uncover various types of changes in (careless) response behavior.
@@
While CODERS is a promising first step towards detecting periods of careless responding rather than identifying which respondents are careless, various pertinent questions remain for future research. For instance, the LSP sequence as a measurement of response invariability may require additional validation, as does the incorporation of response times into CODERS, which we leave for empirical follow-up work. In addition, extending or adapting CODERS with changepoint detection methods that allow for multiple changepoints, or that retain higher power in shorter surveys, may be fruitful directions for further research. Furthermore, it could be interesting to study whether autoencoder reconstruction errors, which can be thought of comparing observed responses with an expectation, are able to distinguish between inconsistent careless responding and other sources of deviations from expected response behavior, such as misunderstanding or misinterpretation of survey items.
@@
CODERS naturally integrates with existing techniques that address careless responding. First, as illustrated in our empirical application, traditional post-hoc screeners and CODERS provide complementary insights. The former aim at identifying which respondents are careless, whereas CODERS specifically looks for changepoints in response behavior. It should be noted that some implementations allow traditional screeners to be computed on subsets of responses. For instance, the implementation of the intra-individual variance (IRV) (Dunn et al.,, 2018; Marjanovic et al.,, 2015) in the R package careless (Yentes & Wilhelm,, 2023) can be calculated for a given number of subsets of equal length. Although this approach is related to the detection of periods of careless responding, it analyzes predefined subsets of responses, while CODERS aims to identify the location of a changepoint.
@@
Second, pre-administered detection items such as bogus, instructed, or self-report items (e.g. Arthur et al.,, 2021; Curran,, 2016; Ward & Meade,, 2023) may provide further validation for the findings of CODERS. An advantage of detection items is that they can in principle detect any type of careless responding. By themselves, detection items do not provide accurate estimates of the location of a change in response behavior, at least when following recommendations of Meade & Craig, (2012) to include at most three detection items, spaced every 50–100 items. However, detection items can strengthen the findings of CODERS: if passed detection items (i.e., responses expected from an attentive participant) fall with the found attentive period while failed detection items fall within the found careless period, there is convergent evidence of such response behavior.
@@
Third, various methods have been proposed in the literature to prevent careless responding in the first place, such as promising rewards for responding accurately (Gibson & Bowling,, 2020), or threatening with punishment for careless responding (Bowling et al.,, 2021). There is evidence that such measures can prevent or at least delay the onset of careless responding (Bowling et al.,, 2021; Gibson & Bowling,, 2020; Huang et al.,, 2012). However, preventive measures can be undesirable or even unethical in practice. For instance, Arthur et al., (2021) and Bowling et al., (2021) advise against using threats if one wishes to maintain a positive long-term relationship with survey participants, while Arthur et al., (2021) stress that promising rewards may cross ethical boundaries in organizational surveys. Notwithstanding, if preventative methods are appropriate, combining them with screening techniques such as CODERS may further enhance quality and reliability of the collected data for subsequent analysis.
@@
Whereas we apply machine learning and nonparametric techniques in CODERS, other literature is focused on model-based approaches. For instance, Ulitzsch et al., 2022b propose a mixture model based on item response theory that is designed to detect careless respondents, which is adapted in Ulitzsch et al., 2022a to incorporate response times. Other approaches that explicitly model carelessness as part of a mixture model are proposed in Arias et al., (2020), Steinmann et al., (2022), and Van Laar & Braeken, (2022). One advantage of model-based approaches is that their estimates can be used to construct person-weights that can be used to downweight the adverse influence of potentially careless respondents in subsequent analyses, which is discussed in detail in Ulitzsch et al., 2022a ; Ulitzsch et al., 2022b and Hong & Cheng, (2019). However, model-based approaches typically impose strong and potentially unverifiable assumptions on the properties of attentive and careless responding.
@@
On a final note, Arthur et al., (2021) envisage “artificial intelligence and machine learning being used to detect careless responding and response distortion in real time in the foreseeable future.” Using machine learning techniques in the form of autoencoders and nonparametric changepoint detection, our proposed method CODERS takes a step towards that future.
@@
Throughout this paper, we have followed recent guidelines by Alfons & Welz, (2024) on the application of machine learning methods for the detection of careless responding. All computations were performed with R version 4.4.2 (R Core Team,, 2024) and Python version 3.10 (Van Rossum & Drake,, 2009). Our implementation of CODERS in the R package carelessonset (Welz & Alfons,, 2024) is available from https://github.com/mwelz/carelessonset. Replication files will be made publicly available upon acceptance of this manuscript and a link will be included here.
@@ Appendix A Setup and Assumptions
This assumption is very mild in survey data since surveys typically measure multiple constructs, with the number of constructs typically known to the survey designer. For instance, the $p=300$ item representation of the NEO-PI-R inventory (Goldberg,, 1999) measures $s=30$ constructs.
@@
This assumption is again mild in survey data. There exist well-established and reliable survey measures for a large variety of constructs. For instance, the International Personality Item Pool (IPIP; Goldberg,, 1999) is a pool of more than 250 personality scales. It is in general recommended to use well-established measures to ensure a highly reliable measurement (Babbie,, 2020, p. 150). The reliability of a measurement can be estimated in multiple ways, such as Cronbach’s alpha, McDonald’s omega and variations thereof, as well as the $H$ coefficient (see, e.g., McNeish,, 2018, for a description of these and other reliability measures). Hence, this assumption is satisfied when the survey data are collected with measures of high reliability.
@@
One way to satisfy this assumption is through randomization of the survey items, which is common practice in surveys in the social and behavioral sciences. If it is desirable to keep items of the same construct together to reduce the burden on respondents, as is often the case in organizational surveys, one may randomize the order of items within each construct, and randomize the order of constructs (or groups of constructs, if related constructs should be presented together). The use of negatively keyed items may also be beneficial.
@@
A detailed motivation for this assumption is provided in the section introducing CODERS.
@@ Appendix B Details on CODERS
B.1 Motivating Novel Methodology
Before we present a technical description of CODERS, we believe that it is instructive to first take a step back and motivate the application of novel techniques for the task of detecting partially careless responding rather than using existing post-hoc screeners such as psychometric antonym, personal reliability, and the longstring index (see e.g., Alfons & Welz,, 2024, for a recent overview of traditional screeners).
A primary motivation for devising new methodology is that most existing screeners compute a single carelessness score for each participant. While this can work very well for distinguishing careless from attentive respondents, having a singular score for every respondent does not provide much information beyond the attentive-careless classification. As outlined in the main text, recent literature suggests that partially careless responding might be a common occurrence. As such, it might be more informative to know when a given respondent starts responding carelessly (if at all) rather than a dichotomous attentive-careless distinction. For instance, this could be useful for knowing when the responses of a given respondent become unreliable so that one must not drop all of their responses in a primary analysis. In order to obtain a granular estimate of carelessness onset on the participant level, we require a carelessness score for every item for every respondent rather than a singular per-participant carelessness score. While some traditional screeners can be somewhat tweaked in this direction (for instance, one could calculate separate intra-individual response variances for separate sets of items; see e.g., Dunn et al.,, 2018; Marjanovic et al.,, 2015), accurate estimation of the onset of carelessness necessitates new methodology.
As such, CODERS is based on two different carelessness scores on the participant-item level to capture a wide variety of different types of careless responding. The first dimension, reconstruction errors (RE) obtained via an auto-associative neural network (autoencoder) (Kramer,, 1992), is intended to measure inconsistent careless responding, such as choosing response categories at random, on the participant-item level. Autoencoders are described in detail in Section B.2. The second dimension, a LongStringPattern (LSP) sequence, is a generalization of the famous longstring index (Johnson,, 2005) that measures on the participant-item level if a given response is part of a content-independent recurring response sequence, such as $-2-1-2-\dots-1-$ 2 - ⋯ - 1 - 2. If so, it allows one to trace back when the recurring sequence has started, thereby providing evidence for the onset of invariable carelessness. LSP is explained in detail in Section B.3.
B.2 Details on Autoencoders
B.2.1 Further Intuition
@@
We stress that response predictions based on a lower-dimensional representation could also be obtained through principal component analysis (PCA). In fact, autoencoders turn out to be a generalization of PCA that, unlike PCA, is not based on linearity assumptions (Kramer,, 1991). Since we wish to refrain from possibly restrictive assumptions, we choose autoencoders over PCA to construct lower-dimensional representations of the given responses. We are now ready to dive into the formal definition of the autoencoder.
B.2.2 Network Architecture and Estimation
@@
Introduced by Kramer, (1992), an autoencoder is a network consisting of an uneven number of $M$ layers, in which the first layer (input layer) holds the input data $\mathbf{x}_{i}$ and the last layer (output layer) holds the reconstructed input data $\widehat{\mathbf{x}}_{i}$. Denoting the number of nodes in the $\ell$-th layer by $N^{(\ell)}$, we have that $N^{(1)}=N^{(M)}=p$. Moreover, the central layer (bottleneck layer) consists of $s<p$ nodes to prevent that the input is simply passed through the network, and the network architecture is symmetric around the bottleneck layer.
@@
for participants $i=1,\dots,n$. The reconstruction of the input vector $\mathbf{x}_{i}$ is then given by $\widehat{\mathbf{x}}_{i}~{}=~{}\bm{f}_{\bm{\theta}}(\mathbf{x}_{i})$. Due to the fact that each node is a function of nodes from a previous layer, the nodes in adjacent layers are often visualized by connecting edges. Figure B.1 provides such a schematic overview of a general autoencoder neural network.
Figure B.1: Schematic overview of an autoencoder neural network with $M$ layers.
@@
$\widehat{\bm{\theta}}=\arg\min_{\bm{\theta}}\Bigg{\{}\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{p}\mathcal{L}\Big{(}x_{ij}-f_{j,\bm{\theta}}(\mathbf{x}_{i})\Big{)}\Bigg{\}}.$ (B.1)
The prediction function of the fitted network is subsequently given by $\bm{f}_{\widehat{\bm{\theta}}}(\cdot)$. For the loss function in (B.1), we choose the smooth and robust Pseudo-Huber loss, which is defined for a fixed $\delta>0$ as
$\mathcal{L}(z)=\delta^{2}\left(\sqrt{1+(z/\delta)^{2}}-1\right).$ (B.2)
This choice results in quadratic loss for small values of $z$ and linear loss for large values of $z$. Consequently, the Pseudo-Huber loss function avoids that large individual reconstruction errors strongly affect the fit, which would make it hard for the network to distinguish attentive responses from inconsistent careless responses, as the latter typically lead to large prediction errors. To solve the optimization problem in (B.1), we use stochastic gradient descent. We refer to Chapter 8.5 in Goodfellow et al., (2016) for details.
B.2.3 Modeling Choices
Following Kramer, (1992), we specify $M=5$ layers in the autoencoder. We set the number of nodes in layers 2 and 4 (mapping an demapping layer, respectively) to $\lfloor 1.5\times p\rfloor$, as a relatively large number of nodes is expected to give the autoencoder flexibility to learn many different types of response behavior (attentive and careless). Based on Assumption 1, the number of nodes in the bottleneck layer equals the number of constructs measured in the questionnaire.
Concerning the activation functions, we again follow a recommendation in Kramer, (1992) and propose to use nonlinear activation functions in the mapping and demapping layers, as well as a linear activation function in the bottleneck layer. Specifically, we propose to use the hyperbolic tangent activation in the (de-)mapping layers and the identity mapping in the bottleneck layer. Definitions are given in Table B.1, which summarizes our proposed autoencoder architecture.
Table B.1: Choices regarding the network architecture of the autoencoder.
@@ B.2.3 Modeling Choices
For fitting the autoencoder, we use the stochastic gradient descent algorithm with a batch size of 10, a learning rate of 0.0001, and 100 epochs. In the pseudo-Huber loss (B.2), we set constant $\delta=1$.
B.3 LongStringPattern Algorithm
@@
Input: Series of observed rating-scale responses $\{x_{j}\}_{j=1}^{p}$, pattern length $l\in\mathbb{N}$ to consider (number of items that an individual pattern encompasses). For instance, the individual pattern $1-2-3$ encompasses $l=3$ items in the response sequence $-3-1-2-$ 3 - 1 - 2 - 3. Note that the algorithm measures straightlining when setting $l=1$.
@@
1
/* fill in the first $l-1$ indices */
2 for $k\leftarrow 1$ to $p-l-1$ by $1$ do
3       if $x[k]==x[k+l]$ then
            increment $streak$ by 1 ;
@@
4
5      else
             Assign value $streak$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{k}$ ;
@@
6             $startindex\leftarrow k+1$;
7             streak $\leftarrow 1$;
8
9       end if
10
11 end for
12
@@ 6             $startindex\leftarrow k+1$;
13 if $\texttt{streak}\geq l$ then
@@
14       Assign value $streak+l-1$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{p}$;
15
16 else if $1<\texttt{streak}<l$ then
@@ 14       Assign value $streak+l-1$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{p}$;
17       Assign value $streak-1$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{p-l}$;
18       Assign value $1$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=p-l+1}^{p}$;
19
20 else
@@ 18       Assign value $1$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=p-l+1}^{p}$;
21       Assign value 1 to all elements in $\{l\textsf{-pattern}_{j}\}_{j=p-l+1}^{p}$;
22
23 end if
@@ 21       Assign value 1 to all elements in $\{l\textsf{-pattern}_{j}\}_{j=p-l+1}^{p}$;
1
/* calculate $l$-pattern for fixed pattern length $l$/
2 for $l\leftarrow 1$ to $l_{max}$ by $1$ do
3       $\{\textsf{LSP}^{l}_{j}\}_{j=1}^{p}\leftarrow l\textsf{-pattern}(\{x_{j}\}_{j=1}^{p},l)$;
4
5 end for
6
@@ 3       $\{\textsf{LSP}^{l}_{j}\}_{j=1}^{p}\leftarrow l\textsf{-pattern}(\{x_{j}\}_{j=1}^{p},l)$;
7 Initialize $\{\textsf{LSP}_{j}\}_{j=1}^{p}$;
8 for $j\leftarrow 1$ to $p$ by $1$ do
9       $\textsf{LSP}_{j}\leftarrow\max\big{\{}\textsf{LSP}_{j}^{1},\textsf{LSP}_{j}^{2},\dots,\textsf{LSP}_{j}^{l}\big{\}}$;
10
11 end for

### v4

@@ Results
Third, there may be participants who alternate between periods of attentive and careless responding (Berry et al., 1992; Clark et al., 2003; Meade and Craig, 2012). Our simulations indicate that CODERS remains conservative in this case and rarely flags one of the changepoints (see Appendix C). CODERS should therefore be viewed as a promising first step towards detecting periods of careless responding, with an extension for the detection of multiple changepoints being an area of further research.
Finally, the changepoint detection procedure of Shao and Zhang (2010) employed by CODERS is based on asymptotic arguments when the number of items is large. Hence, the number of items in the questionnaire is the main determinant for the statistical power of CODERS. For the detection of inconsistent careless responding, the number of participants plays a secondary role regarding power: the more attentive responses the autoencoder can learn from, the better the reconstruction errors can distinguish between attentive and careless responding (cf. Appendix C). The LSP sequence of a given participant, on the other hand, depends only on their own responses, meaning that the number of observations should not play a role concerning power in detecting invariable careless responding. CODERS is thus primarily designed for relatively lengthy questionnaires, and our simulations suggest that there is a gradual drop in statistical power for detecting inconsistency as survey length decreases (see Appendix C). However, partial carelessness due to survey fatigue may be less prevalent in short questionnaires (cf. Bowling et al., 2021). A detailed study of partially careless responding in shorter surveys is therefore beyond the scope of this paper.
@@
In addition to inconsistency and invariability, fast responding is considered the third major manifestation of carelessness (e.g., Ward and Meade, 2023). In (online) surveys, response time is typically measured by the total time a participant spent on the questionnaire, time spent on each questionnaire page, or (less common) time spent on each questionnaire item.
If per-item response times are available, they can easily be incorporated into CODERS as an additional series to be used in the changepoint detection. If only per-page response times are available and pages are relatively short, an additional series can be obtained by assigning to each response the average page time, i.e., the time the participant spent on the page on which the item is located divided by the number of items on that page (which is inspired by the page time index of Bowling et al., 2023). If a participant at some point in the survey, for instance, starts responding without reading the items, we expect a changepoint in response time towards considerably faster responses (cf. Bowling et al., 2023, 2016; Meade and Craig, 2012; Huang et al., 2012).
Nevertheless, it is unclear if a changepoint in response times is indeed indicative of careless responding in an empirical setting. This may depend on the survey design: if items in one part of the survey are sufficiently longer than in other parts, a changepoint in response times might occur naturally for attentive respondents. On the other hand, if all items are of similar length, a changepoint to (much) shorter response times may indicate carelessness. For instance, there exist instruments whose individual items comprise of single words, such as the Big Five unipolar markers (Goldberg, 1992) with items like “timid” or “envious”. In such measurements, one would not expect large differences in response times across items. For such survey designs with similar item length, incorporating response times into CODERS may prove fruitful, but a thorough validation is beyond the scope of this paper and therefore left for future research.
@@ Discussion
We introduced a novel method, CODERS, designed to identify the onset of careless responding (or absence thereof) for each participant in a survey. It is motivated by literature stressing that survey length contributes to careless responding, in the sense that lengthy surveys—which are common in the behavioral sciences—may result in a large proportion of survey participants being partially careless, for instance due to fatigue or boredom. CODERS is highly flexible and can combine multiple indicators for the potential onset of carelessness, such as autoencoder reconstruction errors (RE) as an indicator of inconsistency, and the proposed LongStringPattern (LSP) sequence as an indicator of invariability. Simulation experiments indicate excellent performance of CODERS in identifying the onset of careless responding, and highlight the importance of incorporating multiple indicators that capture different manifestations of carelessness. Re-analyzing data from a seminal study on careless responding, we demonstrate how CODERS reveals novel insights on partial carelessness. In particular when combined with a post-hoc analysis of the identified response segments, CODERS is able to uncover various types of changes in (careless) response behavior.
@@
While CODERS is a promising step towards detecting periods of careless responding rather than identifying which respondents are careless, pertinent questions remain for future research. Most notably, extending or adapting CODERS with changepoint detection methods that allow for multiple changepoints, or that retain higher power in shorter surveys, may be fruitful directions for further research.
@@
CODERS naturally integrates with existing techniques that address careless responding. First, as illustrated in our empirical application, traditional post-hoc screeners and CODERS provide complementary insights. The former aim at identifying which respondents are careless, whereas CODERS specifically looks for changepoints in response behavior. It should be noted that some implementations allow traditional screeners to be computed on subsets of responses. For instance, the implementation of the intra-individual variance (IRV) (Dunn et al., 2018; Marjanovic et al., 2015) in the R package careless (Yentes and Wilhelm, 2023) can be calculated for a given number of subsets of equal length. Although this approach is related to the detection of periods of careless responding, it analyzes predefined subsets of responses, while CODERS aims to identify the location of a changepoint.
@@
Second, pre-administered detection items such as bogus, instructed, or self-report items (e.g. Arthur et al., 2021; Curran, 2016; Ward and Meade, 2023) may provide further validation for the findings of CODERS. An advantage of detection items is that they can in principle detect any type of careless responding. By themselves, detection items do not provide accurate estimates of the location of a change in response behavior, at least when following recommendations of Meade and Craig (2012) to include at most three detection items, spaced every 50–100 items. However, detection items can strengthen the findings of CODERS: if passed detection items (i.e., responses expected from an attentive participant) fall with the found attentive period while failed detection items fall within the found careless period, there is convergent evidence of such response behavior.
@@
Third, various methods have been proposed in the literature to prevent careless responding in the first place, such as promising rewards for responding accurately (Gibson and Bowling, 2020), or threatening with punishment for careless responding (Bowling et al., 2021). There is evidence that such measures can prevent or at least delay the onset of careless responding (Bowling et al., 2021; Gibson and Bowling, 2020; Huang et al., 2012). However, preventive measures can be undesirable or even unethical in practice. For instance, Arthur et al. (2021) and Bowling et al. (2021) advise against using threats if one wishes to maintain a positive long-term relationship with survey participants, while Arthur et al. (2021) stress that promising rewards may cross ethical boundaries in some contexts. Notwithstanding, if preventative methods are appropriate, combining them with screening techniques such as CODERS may further enhance quality and reliability of the collected data for subsequent analysis.
@@
Whereas we apply machine learning and nonparametric techniques in CODERS, other literature is focused on model-based approaches. For instance, Ulitzsch et al. (2022b) propose a mixture model based on item response theory that is designed to detect careless respondents, which is adapted in Ulitzsch et al. (2022a) to incorporate response times. Other approaches that explicitly model carelessness as part of a mixture model are proposed in Zhang et al. (2025), Steinmann et al. (2022), Van Laar and Braeken (2022), and Arias et al. (2020). One advantage of model-based approaches is that their estimates can be used to construct person-weights that can be used to downweight the adverse influence of potentially careless respondents in subsequent analyses, which is discussed in detail in Ulitzsch et al. (2024, 2022a, 2022b) and Hong and Cheng (2019). However, model-based approaches typically impose strong and potentially unverifiable assumptions on the properties of attentive and careless responding.
@@
On a final note, Arthur et al. (2021) envisage “artificial intelligence and machine learning being used to detect careless responding and response distortion in real time in the foreseeable future.” Using machine learning techniques in the form of autoencoders and nonparametric changepoint detection, our proposed method CODERS takes a step towards that future.
@@
[…]
@@ Appendix A Setup and Assumptions
This assumption is very mild in survey data since surveys typically measure multiple constructs, with the number of constructs typically known to the survey designer. For instance, the $p=300$ item representation of the NEO-PI-R inventory (Goldberg, 1999) measures $s=30$ constructs.
@@
This assumption is again mild in survey data. There exist well-established and reliable survey measures for a large variety of constructs. For instance, the International Personality Item Pool (IPIP; Goldberg, 1999) is a pool of more than 250 personality scales. It is in general recommended to use well-established measures to ensure a highly reliable measurement (Babbie, 2020, p. 150). The reliability of a measurement can be estimated in multiple ways, such as Cronbach’s alpha, McDonald’s omega and variations thereof, as well as the $H$ coefficient (see, e.g., McNeish, 2018, for a description of these and other reliability measures). Hence, this assumption is satisfied when the survey data are collected with measures of high reliability.
@@
One way to satisfy this assumption is through randomization of the survey items, which is common practice in surveys in the behavioral sciences. If it is desirable to keep items of the same construct together to reduce the burden on respondents, one may randomize the order of items within each construct, and randomize the order of constructs (or groups of constructs, if related constructs should be presented together). The use of negatively keyed items may also be beneficial.
@@
A detailed motivation for this assumption is provided in the section of the main text that introduces CODERS.
@@ Appendix B Details on CODERS
Motivating Novel Methodology
Before we present a technical description of CODERS, we believe that it is instructive to first take a step back and motivate the application of novel techniques for the task of detecting partially careless responding rather than using existing post-hoc screeners such as personal reliability, psychometric antonym, and the longstring index (see e.g., Alfons and Welz, 2024, for a recent overview of traditional screeners).
A primary motivation for devising new methodology is that most existing screeners compute a single carelessness score for each participant. While this can work very well for distinguishing careless from attentive respondents, having a singular score for every respondent does not provide much information beyond the attentive-careless classification. As outlined in the main text, recent literature suggests that partially careless responding might be a common occurrence. As such, it might be more informative to know when a given respondent starts responding carelessly (if at all) rather than a dichotomous attentive-careless distinction. For instance, this could be useful for knowing when the responses of a given respondent become unreliable so that one must not drop all of their responses in a primary analysis. In order to obtain a granular estimate of carelessness onset on the participant level, we require a carelessness score for every item for every respondent rather than a singular per-participant carelessness score. While some traditional screeners can be somewhat tweaked in this direction (for instance, one could calculate separate intra-individual response variances for separate sets of items; see e.g., Dunn et al., 2018; Marjanovic et al., 2015), accurate estimation of the onset of carelessness necessitates new methodology.
As such, CODERS is based on two different carelessness scores on the participant-item level to capture a wide variety of different types of careless responding. The first dimension, reconstruction errors (RE) obtained via an auto-associative neural network (autoencoder) (Kramer, 1992), is intended to measure inconsistent careless responding, such as choosing response categories at random, on the participant-item level. The second dimension, a LongStringPattern (LSP) sequence, is a generalization of the famous longstring index (Johnson, 2005) that measures on the participant-item level if a given response is part of a content-independent recurring response sequence, such as $1-2-1-2-\dots-1-2$. If so, it allows one to trace back when the recurring sequence has started, thereby providing evidence for the onset of invariable carelessness. Autoencoders and LSP are described in detail in the following sections.
Details on Autoencoders
Further Intuition
@@
We stress that response predictions based on a lower-dimensional representation could also be obtained through principal component analysis (PCA). In fact, autoencoders turn out to be a generalization of PCA that, unlike PCA, is not based on linearity assumptions (Kramer, 1991). Since we wish to refrain from possibly restrictive assumptions, we choose autoencoders over PCA to construct lower-dimensional representations of the given responses. We are now ready to dive into the formal definition of the autoencoder.
Network Architecture and Estimation
@@
Introduced by Kramer (1992), an autoencoder is a network consisting of an uneven number of $M$ layers, in which the first layer (input layer) holds the input data $\mathbf{x}_{i}$ and the last layer (output layer) holds the reconstructed input data $\widehat{\mathbf{x}}_{i}$. Denoting the number of nodes in the $\ell$-th layer by $N^{(\ell)}$, we have that $N^{(1)}=N^{(M)}=p$. Moreover, the central layer (bottleneck layer) consists of $s<p$ nodes to prevent that the input is simply passed through the network, and the network architecture is symmetric around the bottleneck layer.
@@
for participants $i=1,\dots,n$. The reconstruction of the input vector $\mathbf{x}_{i}$ is then given by $\widehat{\mathbf{x}}_{i}~=~\bm{f}_{\bm{\theta}}(\mathbf{x}_{i})$. Due to the fact that each node is a function of nodes from a previous layer, the nodes in adjacent layers are often visualized by connecting edges. Figure B1 provides such a schematic overview of a general autoencoder neural network.
Figure B1: Schematic overview of an autoencoder neural network with $M$ layers.
@@
$\widehat{\bm{\theta}}=\arg\min_{\bm{\theta}}\Bigg\{\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{p}\mathcal{L}\Big(x_{ij}-f_{j,\bm{\theta}}(\mathbf{x}_{i})\Big)\Bigg\}.$ (B1)
[…]
$\mathcal{L}(z)=\delta^{2}\left(\sqrt{1+(z/\delta)^{2}}-1\right).$ (B2)
This choice results in quadratic loss for small values of $z$ and linear loss for large values of $z$. Consequently, the Pseudo-Huber loss function avoids that large individual reconstruction errors strongly affect the fit, which would make it hard for the network to distinguish attentive responses from inconsistent careless responses, as the latter typically lead to large prediction errors. To solve the optimization problem in (B1), we use stochastic gradient descent. We refer to Chapter 8.5 in Goodfellow et al. (2016) for details.
Modeling Choices
Following Kramer (1992), we specify $M=5$ layers in the autoencoder. We set the number of nodes in layers 2 and 4 (mapping an demapping layer, respectively) to $\lfloor 1.5\times p\rfloor$, as a relatively large number of nodes is expected to give the autoencoder flexibility to learn many different types of response behavior (attentive and careless). Based on Assumption 1, the number of nodes in the bottleneck layer equals the number of constructs measured in the questionnaire.
Concerning the activation functions, we again follow a recommendation in Kramer (1992) and propose to use nonlinear activation functions in the mapping and demapping layers, as well as a linear activation function in the bottleneck layer. Specifically, we propose to use the hyperbolic tangent activation in the (de-)mapping layers and the identity mapping in the bottleneck layer. Definitions are given in Table B1, which summarizes our proposed autoencoder architecture.
Table B1: Choices regarding the network architecture of the autoencoder.
@@
For fitting the autoencoder, we use the stochastic gradient descent algorithm with a batch size of 10, a learning rate of 0.0001, and 100 epochs. In the pseudo-Huber loss (B2), we set constant $\delta=1$.
LongStringPattern Algorithm
@@
Input: Series of observed rating-scale responses $\{x_{j}\}_{j=1}^{p}$, pattern length $l\in\mathbb{N}$ to consider (number of items that an individual pattern encompasses). For instance, the individual pattern 1231-2-3 encompasses $l=3$ items in the response sequence 1231231-2-3-1-2-3. Note that the algorithm measures straightlining when setting $l=1$.
@@
/* fill in the first l1l-1 indices */
for $k\leftarrow 1$ to pl1p-l-1 by 11 do
 if $x[k]==x[k+l]$ then
     increment streakstreak by 1 ;
@@
 else
      Assign value streakstreak to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{k}$ ;
@@
    $startindex\leftarrow k+1$;
      streak $\leftarrow 1$;
   end if
end for
@@
if $\texttt{streak}\geq l$ then
@@
   Assign value streak+l1streak+l-1 to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{p}$;
else if $1<\texttt{streak}<l$ then
@@
   Assign value streak1streak-1 to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{p-l}$;
   Assign value 11 to all elements in $\{l\textsf{-pattern}_{j}\}_{j=p-l+1}^{p}$;
else
@@
   Assign value 1 to all elements in $\{l\textsf{-pattern}_{j}\}_{j=p-l+1}^{p}$;
end if
@@
/* calculate $l$-pattern for fixed pattern length $l$ */
for $l\leftarrow 1$ to $l_{max}$ by 11 do
 $\{\textsf{LSP}^{l}_{j}\}_{j=1}^{p}\leftarrow l\textsf{-pattern}(\{x_{j}\}_{j=1}^{p},l)$;
end for
@@
Initialize $\{\textsf{LSP}_{j}\}_{j=1}^{p}$;
for $j\leftarrow 1$ to $p$ by 11 do
 $\textsf{LSP}_{j}\leftarrow\max\big\{\textsf{LSP}_{j}^{1},\textsf{LSP}_{j}^{2},\dots,\textsf{LSP}_{j}^{l}\big\}$;
end for